4x-y=-1
6x+5y=18
I really want to know the complete solution, and how to do it. May god bless the one who answers me.

Found 2 solutions by borzeya, kiru_khandelwal:
Answer by borzeya(6)

(-6)*(4x-y=-1)
(4)*(6x+5y=18)
-24x+6y=6
24x+20y=72
26y=78
y=3
4x-y=-1
4x-3=-1
4x=2
x=0.5

Answer by kiru_khandelwal(79)

4x -y = -1
=> 4x + 1 = y
=> y = 4x + 1........Equation (1)
6x + 5y = 18...equation(2)
Substituting the value of y from equation(1) in the equation(2)
=> 6x + 5(4x+1) = 18
=> 6x + 20x + 5 = 18
=> 26x = 18 -5
=> 26x = 13
=> x = 13/26 = 1/2
substituting the value of x in equation(1)
y = 4x + 1
y = 4*1/2 + 1
y = 2+1 = 3
So, x=1/2 y=3
Hope this helps :-)